# Hydra: Virtualized Multi-Language Runtime for High-Density Serverless Platforms

### v2

@@ 3.1. Function Interface
This function interface is heavily inspired by existing function interfaces available in both commercial (Amazon Lambda (lambda, ), Azure Functions (functions, ), and Google Cloud Functions (gcf, )), and open-source Serverless platforms (such as OpenFaaS (openfaas, ), OpenWhisk (openwhisk, ), kative (knative, )). We expect that Hydra requires minor to no function source code modifications to run existing serverless functions in Hydra.
3.2. Function Sandboxes
Hydra sandboxes are based on GraalVM’s Native Image (wimmer:2019, ) isolates which we extended to support on-demand sandbox creation, payload copying, function invocation, and checkpoint/restore. Native Image is an Ahead-Of-Time Java compiler that we use to compile user functions into shared libraries. During the Native Image compilation process (from Java bytecode into a shared native library), we introduce compiler extensions that inject additional Application Binary Interface (ABI) functions into the final shared library.
Upon receiving a request to invoke a function, Hydra loads the function shared library using dlopen, if not already loaded. The shared library loads into memory the function code cache and the initial heap state (see Figure 4). The code cache contains the compiled user code that will be invoked through the registered entrypoint and the initial heap state contains a set of global values which are copied into the isolate memory heap upon modification (wimmer:2019, ) (copy-on-write). Both the code cache and the initial heap state are shared among all memory isolate sandboxes of a single function. Different functions have separate code caches and isolate heap states.
Listing 3.1 shows the most relevant functions exposed by shared libraries. The create_isolate creates an isolate sandbox and returns a pointer to the new isolate (through the second function argument). Once an isolate is created, threads can be attached to it through the attach_thread function. This function creates a new thread stack in the given isolate and returns a pointer to the thread data structure (which can be later used to detach the thread from the isolate by calling detach_thread). Note that threads can be created in any isolate and will start executing in the same isolate as the parent thread. To copy the invocation arguments and return values, we use the copy function that copies a serialized buffer of characters into a destination isolate and returns a handle to it. This handle can later be used to retrieve the copied values once inside the destination isolate.
Functions are invoked through invoke_function, passing the target isolate where the request will be handled, the isolate where the result value should be copied into (Hydra isolate), the entrypoint, and arguments handle (previously copied with copy). The function returns a handle that can be used to retrieve the result value back in Hydra.
@@ 3.3. Multi-Language Support in Hydra
To run functions written in languages other than Java, Hydra takes advantage of Truffle (wuerthinger:2017, ), a Java-based language implementation framework. We extend the Truffle framework to i) transparently load and configure the correct Truffle language, ii) load the function code into the Truffle engine, and iii) integrate it with Hydra’s sandbox ABI (introduced in Listing 3.1). We AOT compile the Truffle language implementation bundled together with the function code into a dynamic library that is then registered into Hydra. The resulting library supports Hydra’s sandbox ABI. The function entry point jumps to a call gate injected by us that enforces the correct translation of Java types into the target language types (used to pass invocation arguments and return values).
@@
Truffle language implementations are interpreted and Just-In-Time (JIT) compiled. As a consequence, the code cache will be updated by the Truffle JIT compile that comes built-in the shared library. Since multiple isolate sandboxes of the same function share the code cache, the optimized JIT code will be shared across all active isolate sandboxes for functions written in Truffle languages.
3.4. Function Snapshotting
Function sandboxes will become bloated when JIT-compiled languages are utilized (via Truffle). To reduce the startup time and memory footprint of these sandboxes, we propose snapshotting function sandboxes after the first initial requests (for example, when a sandbox is to be shut down due to inactivity). The snapshot needs to include not only the memory heap but also the code cache and initial heap state. This way, all the code interpretation and JIT compilation can be bypassed when the sandbox is restored.
@@ 3.4. Function Snapshotting
Snapshotting a single sandbox is not trivial as it requires saving all memory regions and resources that pertain to a single sandbox. Taking CRIU (criu, ) as a reference, C/R is done at the process level by inspecting the entire set of resources utilized by a process and dumping them to one or more files to be later used to restore the entire process. In Hydra, this is not possible as we want to snapshot a single sandbox, and it is not possible to determine which memory segments and files belong to a particular sandbox at checkpoint time. Moreover, Hydra does not control where and how the sandbox memory heap grows, nor where the code cache lives, as this would require invasive changes to the Native Image and Truffle.
To overcome this challenge, Hydra tracks the execution of a function invocation deemed for checkpoint. This is done by intercepting system calls and tracking their output. By compiling all the resources utilized by the sandbox, it is possible to dump its state and restore it later. Our current implementation focuses on tracking memory mappings and open files, which we found to be sufficient to support functions in Hydra. Nevertheless, the proposed design is generic and can be easily extended to monitor more system calls and checkpoint/restore resources other resources.
@@
3.4.1. Tracing and Dumping a Function
Upon checkpointing a function, Hydra attaches a seccomp-bpf (seccomp, ) filter to the thread handling the function invocation. After installing the filter, the thread loads the function library (dlopen), creates an isolate, and invokes the function code. During this entire process, another thread stays behind receiving system call notifications. We monitor system calls related to memory (such as mmap, munmap, and mprotect) and files (such as openat, dup, and dup).
Upon receiving a system call notification, the monitor thread registers the arguments and the return value. This information is kept in a list in memory. In addition to keeping track of the received system calls, the monitor thread also keeps in memory the current state of memory mappings and open file descriptors belonging to the target sandbox.
For memory-related operations, the monitor keeps a list of memory mappings including not only the range but also their current permission and if the memory in that mapping was ever writable. We keep this information at the page granularity level. A similar table is kept for file-related operations, tracking which file descriptors are in use.
The monitor thread waits until the monitored thread exists the function code and detaches from the isolate to proceed with the checkpoint operation. At this moment, the monitor starts by iterating over the list of system calls to remove unnecessary entries. For example, a file may be opened, its contents read into memory, and then closed. In this case, we do not need to keep the open and close system calls as we are only trying to restore the final stage of execution of the function code. The same principle can be applied to memory operations where mmap system calls whose return address was already unmmaped can be safely ignored. The final list of system calls (including arguments and return values) is then written to a file used during the restore operation.
The second checkpoint step is to iterate over all memory mappings and copy the memory contents to a file. We can also optimize this process by copying only the mappings that could have been modified (i.e., write permission was enabled at the point in time). Note that all the read-only memory will be restored when the system calls that loaded that memory are replayed. Two files result from a checkpoint operation: i) a file containing the set of system calls that need to be replayed upon a restore, and ii) the memory contents that need to be restored.
3.4.2. Restoring a Function from a Snapshot
Hydra can restore a function sandbox (including the code cache and the initial heap state) from the snapshot produced during a checkpoint operation. Hydra starts by loading the file containing the system calls that need to be replayed. For each system call, the restoring thread replays the system call and confirms that the output is the one expected. Note that for mmap system calls we force address to be the same as the one returned by the original mmap system call by passing the target address and the MAP_FIXED flag. After restoring all system calls, memory contents and open files should match the state of the original state.
@@ 3.4.2. Restoring a Function from a Snapshot
After this initial step, the restoring thread mmaps the memory mappings file contents to the expected memory positions. This technique significantly speeds up the restoring operation at the expense of some additional page faults during the first invocation of the function (note that this could be further optimized by pre-fetching the pages estimated to be necessary as done in previous work (ustiugov:2021, ).
3.4.3. Supporting Multiple Functions
@@
The C/R mechanism described so far can easily fail if two function snapshots happen to require the same virtual address range. This can happen, for example, if two calls to mmap return the same address in two different executions of Hydra where each function was individually checkpointed. Without a reliable mechanism to address this issue, restore operations can easily fail if the memory is already occupied.
@@
To prevent this scenario, Hydra controls which memory ranges are given to each sandbox during the checkpoint phase. This is possible by manipulating the mmap arguments and forcing each function to operate on disjoint memory ranges. Note that this mechanism works as most runtimes reserve large blocks of memory which are then used internally to feed allocation pools (e.g., memory heaps).
@@
Hydra reserves 1 TB of virtual memory for each function. This should be enough to accommodate the memory needs of any serverless function and, given a 64-bit address space, should also not pose any restrictions on the number of functions that can be loaded at the same time. We use a seed value that is unique for each function and that is used to pick a 1 TB range that does not overlap with any other function’s virtual memory range. Moreover, this mechanism can be extended to create multiple snapshots of the same function using different virtual memory ranges as a way to reintroduce ASLR. This is, when restoring a function snapshot, the platform could randomly decide which snapshot to use to load the same function into different memory ranges.
3.5. Forked Function Sandbox
@@
Hydra also supports creating sandboxes with stronger isolation by handling function invocations in sub-processes. Such sandboxes are necessary to isolate functions that execute unmanaged code, i.e., code that executes outside the sandbox and that may not be ready to receive multiple concurrent invocations. For example, multiple functions executing managed code (Python code for example) can have their execution properly contained using memory isolates as the runtime controls in which memory heap objects are allocated, which threads belong to which isolate, etc. If a particular thread invokes a native method, the isolation guarantees enforced by the runtime no longer apply and the thread may be able to access memory that belongs to other isolates. Furthermore, external libraries may also contain state that is not properly isolated across concurrent function invocations. Deciding if a function should execute in a forked sandbox is outside the scope of this work but could be done by checking if a function includes native dependencies at compile or deployment time.
@@
Functions ran in forked sandboxes are initially loaded into the parent Hydra process. This ensures that both the code cache and the initial heap state are shared in a copy-on-write manner with all sandboxes deployed in child processes. Upon creating a new sandbox, Hydra invokes the Native Sandbox Interface (NSI in Figure 4); if process-isolation is enabled, a forked process is created. To communicate between the parent and child processes we use two pipes (returned by pipe): one to send invocation requests from the parent to the child, and one to receive the invocation result from the child to the parent. Both pipes are created right before forking. Upon creating a new Fork sandbox, the child process closes parent file descriptors (e.g., the file descriptor used to receive REST requests) and resets signal handlers. Then it creates a new sandbox isolate and finally blocks the reading pipe waiting for arguments for a function invocation. The parent side simply creates a process handle with references to both pipes.
@@ 4. High-Density Serverless Platform
In addition to Hydra, we also built a simple serverless platform to evaluate how co-location could affect scheduling policies and deployment density. The top-level component of our platform is the scheduler, which acts as an entry point for the platform. The scheduler controls worker nodes running Node Managers, which in turn manage Hydra instances capable of executing the actual function code. Figure 6 shows the high-level architecture of the serverless platform.
4.1. Node Manager
The node manager governs Hydra instances in a local node. It is responsible for managing the lifetime of Hydra instances, scheduling invocations across those instances, and forwarding invocation requests to them. When scheduling invocations, the node manager attempts to prioritize co-locating invocations coming from the same tenant in a single Hydra instance to benefit from sharing runtime components and avoiding cold starts. If the instance cannot accommodate another incoming invocation, a new instance is picked to serve this invocation. After invocation execution finishes, the node manager keeps the Hydra instance alive for a configurable amount of time to maintain it warm for a potential subsequent invocation.

### v3

@@ 3.2. Function Sandboxes
Functions are invoked through invoke_sandbox, passing the target sandboxed thread handle where the request will be handled, a buffer where the arguments should be read from (for example, a JSON represented as a C string), and another buffer where result should be placed (also as a C string). The final argument indicates the maximum output size. All functions return an error code.
@@ 3.3. Multi-Language Support in Hydra
Native Image only supports Java functions. To run functions written in languages other than Java, Hydra takes advantage of Truffle (wuerthinger:2017, ), a Java-based language implementation framework. We extend the Truffle framework to i) transparently load and configure the correct Truffle language, ii) load the function code into the Truffle engine, and iii) integrate it with Hydra’s sandbox ABI (introduced in Listing 3.2). Using Native Image, we AOT compile the Truffle language implementation bundled together with the function code into a dynamic library that is then registered into Hydra. The resulting library supports Hydra’s sandbox ABI. The function entry point jumps to a call gate injected by us that enforces the correct translation of Java types into the target language types (used to pass invocation arguments and return values).
@@
Truffle language implementations are interpreted and Just-In-Time (JIT) compiled. As a consequence, the code cache will be updated by the Truffle JIT compile that comes built-in the shared library. Multiple sandboxes running same function share the code cache this reusing already profiled and optimized code.
3.4. Hydra Sandbox Snapshots
Function sandboxes require significant warmup time and memory, especially for JIT-compiled languages such as JavaScript. To reduce the startup time and memory footprint of these sandboxes, Hydra is capable checkpointing function sandboxes. A snapshot includes the entire sandbox state such that all the code interpretation and JIT compilation overheads can be bypassed upon restore.
@@ 3.4. Hydra Sandbox Snapshots
Checkpointing a single sandbox is not trivial as it requires saving all memory regions and resources that pertain to a single sandbox. Taking CRIU (criu, ) as a reference, C/R is done at the process level by inspecting the entire set of resources utilized by a process and dumping them into one or more files to be later used to restore the entire process. In Hydra, this is not possible as we need to checkpoint a single sandbox and not the entire process. For instance, it is not possible to determine which memory mappings and file descriptors belong to a particular sandbox as the OS organizes resources by process.
To overcome this challenge, Hydra tracks the execution of a sandbox by intercepting system calls and tracking their side effects. By keeping track of all the resources utilized by a sandbox, it is possible to checkpoint its state and restore it later. Hydra’s sandbox checkpoint/restore design is focused on four main elements we found to be crucial for Java, Python, and JavaScript applications: memory mappings (used by language engines to allocate managed heaps), file descriptors, libc memory allocator (essential to many libc functions used by language engines), and threads running in the sandbox.
3.4.1. Function Sandbox Checkpoint
Hydra utilizes seccomp-bpf to selectively intercept system calls by attaching a seccomp-bpf (seccomp, ) filter to the thread handling the function invocation. After installing the filter, the thread loads the function library with dlopen, creates a sandbox, and invokes the function code. Another thread is used track system call notifications and monitors system calls related to memory (such as mmap, munmap, and mprotect), files (such as openat, dup), and threads (clone, clone3, exit). We will focus on these three main classes of system calls and omit, for simplicity, other less important system calls and system calls variants.
Upon receiving a system call notification, the monitor thread registers the arguments and the return value. This information is kept in a data structure in memory. In addition to keeping track of the received system calls, the monitor thread also keeps the current state of memory mappings, open file descriptors, and running threads belonging to the target sandbox.
Memory mappings and file descriptors. For memory-related system calls, the monitor keeps a list of memory mappings including not only the range but also their current permission and if the memory in that mapping was ever writable. Hydra keeps this information at the page granularity. A similar table is kept for file-related operations, tracking which file descriptors are in use.
The monitor thread waits until the function invocation finishes and detaches the invoking thread from the sandbox to proceed with the checkpoint operation. At this point, the monitor starts iterating over the list of system calls to remove unnecessary entries. For example, a file may be opened, its contents read into memory, and then closed. In this case, Hydra discards these open and close system calls as it targets restoring the final state sandbox. The same principle is applied to memory operations where mmap system calls whose return address was already unmmaped can be safely ignored. The final list of system calls (including arguments and return values) is then written to a file used during the restore operation. For the benchmarks used in section 5, the final number of system calls was reduced by 47.1%. on average.
The second checkpoint step is to iterate over all memory mappings and copy the memory contents to a file. We optimize this stage by copying only the mappings that could have been modified (i.e., write permission was enabled at some point in time). Note that all read-only memory will be restored when the system calls that loaded that memory are replayed (for instance, an mmap that loads a binary into memory). Two files result from a checkpoint operation: i) a file containing the set of system calls that need to be replayed upon a restore, and ii) the memory contents that need to be restored.
Threads. Similarly to memory and file descriptors, Hydra keeps track of thread creation and destruction by monitoring system calls. However, saving the state of a thread requires stopping it and saving its CPU state. In Hydra, we do this by freezing all threads associated with the sandbox being checkpointed (this association is kept by monitoring system calls) and sending them a signal (we use SIGUSR1 but other signals could be used). Background threads running in the sandbox are frozen while the checkpoint operation is in progress, and are resumed after, allowing the sandbox to receive further invocations.
After receiving a signal, and once inside the signal handler installed by Hydra, Hydra saves the ucontext_t passed as argument into the handler. This data structure contains the CPU state of the thread (general-purpose register values, stack pointer, instruction pointer, etc.) necessary to restore the thread later on. Note that the thread state is saved while the thread is inside the signal handler as this will be important for the restore operation.
Memory allocator. Besides memory mappings resulting from mmap, libc’s memory allocator, malloc, may also be called to allocate memory. For instance, creating a thread with pthread_create will allocate memory through malloc. Since multiple sandboxes share the same underlying libc, it becomes important to separate the memory chunks allocated with libc’s allocator for different sandboxes.
To separate the allocated memory for different sandboxes, Hydra offers a virtualized memory allocator that overrides libc’s memory allocation interface (malloc, free, etc.). Hydra’s memory allocator is based on the Doug Lea’s memory allocator (malloc, ), which includes support for memory spaces. Hydra’s memory allocator uses one memory space per sandbox and supports checkpoint and restore operations. Upon a call to the allocator (e.g., a call to malloc) the allocator starts by checking if the thread already belongs to a memory space. This check is based on a thread-local variable (fast-path), or by checking a table kept in the allocator metadata (slow-path). When a checkpoint operation is requested, the allocator writes to disk all of its internal metadata regarding a particular sandbox.
3.4.2. Function Sandbox Restore
Hydra can restore a function sandbox (including the code cache and the initial heap state) from the snapshot produced during a checkpoint operation. Hydra starts by loading the file that contains the system calls that need to be replayed. For each system call, the restoring thread replays the system call and confirms that the output is the one expected (e.g., checking that the file descriptor number is the same as the one obtained in the original execution before the checkpoint). Note that for mmap system calls, Hydra forces the address to be the same as the one returned by the original mmap system call by passing the target address and the MAP_FIXED flag. After replaying all memory map and file-related system calls, memory contents and open files match the state of the original sandbox.
@@ 3.4.2. Function Sandbox Restore
After this initial step that restores file descriptors, file-based memory mappings, and read-only memory mappings, the restoring thread loads the memory.dump file contents to the expected memory positions (see Figure 6). This file includes only memory mappings recorded in the original execution that may have been dirtied. This technique significantly speeds up the restore operation at the expense of some additional page faults during the first invocation of the function (note that this could be further optimized by prefetching the pages estimated to be necessary as done in previous work (ustiugov:2021, )).
In the third step, Hydra restores the memory allocator state by loading the memory space metadata into Hydra’s memory allocator. And finally, Hydra restores sandbox threads. Thread restoration is achieved with the help of the sigreturn system call that receives the ucontext_t structure passed into the signal handler in the checkpoint step. The sigreturn system call allows the kernel to restore the state of the thread by setting all its registers. The thread is restored in its original application as if it had just received a signal. System calls interrupted during the checkpoint step (e.g. futex) are automatically re-started upon a restore. Finally, we ensure that threads are restored using the same thread id by controlling the value in ns_last_pid file in the proc file system.
3.4.3. Restoring Multiple Snapshots
@@
The C/R mechanism described so far can easily fail if two function snapshots require the same virtual address range. This can happen, for example, if two calls to mmap return the same address in two different executions of Hydra where each function was individually checkpointed. Without a reliable mechanism to address this issue, restore operations can easily fail if the virtual memory is already being used.
@@
To avoid this scenario, Hydra controls which memory ranges are assigned to each sandbox. This is possible by intercepting and adjusting the mmap arguments and forcing each function to operate in disjoint memory ranges. Note that this mechanism works as most language engines reserve large blocks of memory which are then used internally to feed allocation pools (e.g., memory heaps). Hydra’s memory allocator also relies on mmap to serve memory arenas.
@@
Hydra reserves 32 GBs of virtual memory for each function, although this value is configurable. We use a unique seed value that is used to pick a 32 GB range that does not overlap with any other sandboxes’ virtual memory range. Moreover, this mechanism allows the creation of multiple snapshots of the same function using different virtual memory ranges as a way to reintroduce Address space layout randomization (ASLR). This way, when restoring a function snapshot, the platform could randomly decide which snapshot to use to load the same function into different memory ranges.
The same seed value is also used separate file descriptor and thread id ranges between multiple snapshots. In particular, this ensures that a after restoring a snapshot, the function code will not mistakenly use a file descriptors from another function.
3.4.4. Ignored system calls and limitations.
Hydra records and replays system calls that modify system resources. System calls that do not modify the system state such as getpid and sched_yield are not monitored. Other system calls that change the state of the system are either recorded (mmap, open, clone, etc), or its state is reproduced. For example, read and write are not recorded, but advances in a file-based file descriptor offset can be reproduced by moving the file offset upon a restore operation.
Hydra’s snapshots currently do not try to repair network connections. From our experience, language engines quickly reconnect with remote endpoints after a restore. Our current prototype currently also does not checkpoint sub-processes, which could be checkpointed by integrating CRIU (criu, ). We have not found examples of serverless applications that left background processes after an invocation. We assume that files produced by the function code will be transported to the restore destination and will not have been modified.
3.5. Functions with Native Extensions
@@
Serverless functions often invoke native extensions in the form of native libraries. These native extensions, commonly used in machine learning inference and video processing workloads for example, extensively use advanced hardware features (such as SIMD and cryptographic instructions) that are not supported by the managed language runtime where the function is running. Since native extensions run outside memory isolate sandboxes, where SFI no longer applies, Hydra supports creating sandboxes with stronger isolation by handling function invocations in sub-processes. Such sandboxes are necessary to isolate functions that execute native code, i.e., code that executes outside the sandbox and that may not be ready to receive multiple concurrent invocations. For instance, multiple function invocations executing Python code can have their execution properly contained using sandboxes as the language engine controls in which memory heap objects are allocated. If a particular thread invokes a native method, the isolation guarantees enforced by the engine no longer apply and the thread may be able to access memory that belongs to other functions or corrupt its own managed memory. Furthermore, native libraries may also contain state that is not properly isolated across concurrent function invocations. Hydra can also restore function snapshots with native extensions in a sub-process, allowing functions to benefit from a faster cold start by skipping the initialization of the sandbox entirely in the sub-process. Deciding if a function should execute in a forked sandbox is outside the scope of this work, but could be achieved by analyzing the function code to identify native dependencies at compile or deployment time.
@@
Functions running in sub-process sandboxes are initially loaded into the parent Hydra process. This ensures that both the code cache and the initial heap state are shared in a copy-on-write manner with all sandboxes deployed in child processes. Upon creating a new sandbox, Hydra invokes the Native Sandbox Interface (NSI in Figure 5); if process-isolation is enabled, a sub-process is created. To communicate between the parent and child processes Hydra uses two pipes (returned by the pipe system call): one to send invocation requests from the parent to the child, and one to receive the invocation result from the child to the parent. Both pipes are created right before forking. Upon creating a new forked sandbox, the child process closes parent file descriptors (e.g., the file descriptor used to receive REST requests) and resets signal handlers. Then it creates a new memory sandbox and finally blocks the reading pipe waiting for arguments for a function invocation. The parent side simply creates a process handle with references to both pipes.
@@ 4. High-Density Serverless Platform
In addition to the proposed runtime, Hydra features a surrounding serverless platform with colocation-aware scheduling policies to maximize deployment density. The top-level component of our platform is the scheduler, which acts as an entry point for the platform. The scheduler controls nodes running Hydra Managers, which in turn manage Hydra runtimes capable of executing the actual function code. Hydra managers maintain a pool of fresh Hydra instances ready to host any function to avoid initializing the virtualization stack on the cold start critical path. Figure 7 shows the high-level architecture of the serverless platform.
4.1. Hydra Scheduler
@@
To fully leverage Hydra runtime colocation capabilities, we propose a scheduler featuring additional policies to maximize function density. The scheduler governs a fixed number of nodes (Hydra Managers) and keeps track of the state of each node to make scheduling decisions and balance the load. The state of a node includes the node’s registered tenants and functions, and estimated memory utilization. In Hydra, functions can have different sizes, determining the maximum memory that can be allocated for this function’s sandbox. Hydra scheduler estimates the memory utilization of the node by aggregating the sizes of the functions currently running invocations. Based on the state of the nodes, the scheduler attempts to group invocations in the same nodes (node capacity permitting) using functions and then tenants as a criterion for colocation. Finally, if none of the nodes already have the function being invoked or any other function from the same tenant registered, the scheduler selects the least utilized node to serve the invocation.
The scheduler supports four colocation modes: i) full colocation (all invocations from all tenants can be colocated in the same Hydra runtime); ii) tenant-based colocation (only functions of the same tenant can be colocated, default for Hydra); iii) single-function colocation (only invocations of the same function can be colocated); iv) and no colocation, the execution mode offered by most commercial and open-source serverless platforms. Existing schedulers have no incentive to explore modes i) and ii) because state-of-the-art runtimes do not colocate different functions within a single runtime. Runtimes like Knative (knative, ) mostly benefit from mode iii) since Knative supports concurrent invocations to the same function within a single runtime. Mode iv) is suitable for OpenWhisk (openwhisk, ) since OpenWhisk runtimes do not support colocation.
4.2. Hydra Manager
The Hydra Manager runs Hydra instances in a local node. It is responsible for managing the lifetime of Hydra instances and scheduling invocations across those instances. When scheduling invocations, the Hydra Manager prioritizes colocating them in a single Hydra instance to benefit from sharing runtime components and avoiding cold starts according to the colocation mode configured in the upstream scheduler, described in section 4.1. Hydra instances running invocations in a node are flexible, i.e., they can grow to accommodate parallel invocations and shrink to return memory to the OS. After invocation execution finishes, the Hydra Manager keeps the Hydra instance alive for a configurable amount of time to keep it warm for a potential subsequent invocation.
Hydra runtime is function- and language-generic, enabling Hydra platform to maintain a pool of Hydra instances capable of hosting any function. Instead of creating instances on demand, Hydra Manager fetches a pre-created instance from this pool. This approach avoids initializing the Hydra instance and VM/container on the cold start critical path. As discussed in section 2.3, with runtimes like OpenWhisk (openwhisk, ), platforms can also maintain a pool of instances. However, OpenWhisk runtime instances are language-specific, implying that platforms would have to maintain separate pools for each language, whereas the Hydra platform only needs to maintain a single pool without speculatively partitioning it between languages. For Knative, platforms cannot pre-create instances since Knative instances are function-specific. Besides, the Hydra platform does not have to resize Hydra instances when registering a function since Hydra runtime can create memory-constrained sandboxes (factor #2 in section 2.3). In addition, Hydra runtime can destroy sandboxes after a configurable period of inactivity, effectively returning memory to the OS, unlike Knative, where concurrent requests share the same heap that cannot shrink (factor #1 in section 2.3).